# Effect of Strain on the Band Gap of Monolayer MoS2

### v2

@@ V Acknowledgements
The work of RKS was supported by the Department of Energy, Office of Science, Basic Energy Sciences, Materials Chemistry Materials Science and Engineering, under grant DE-DC0023356. The work of RKS and HT was partially supported by the Temple Center for Research on Energy Materials through the Catalytic Collaborative Funding Initiative sponsored by the Office of the Vice President for Research at Temple. The work of JPP was supported by DOE grant DE-DC0023356, and by the National Science Foundation, Division of Materials Research, Condensed Matter and Materials Theory, under grant DMR-2344734. We thank Professor Maria Iavarone for helpful discussions and suggestions.
VI Supplementary Information
Possible explanation for the decrease in band gap with strain.
@@ VI Supplementary Information
In the 1H MoS2 monolayer, one Mo atom and its six adjacent S atoms form a trigonal prism (see Figure S1), and in the ligand field theory, the d-orbitals of the transition metal atoms split into three groups of non-bonding levels Chhowalla et al. (2013). The lower split group is mainly of $d_{z^{2}}$ character and is filled, and the fundamental gap is formed by this filled group and the unoccupied middle group.
To understand the trend in the change of gap with applied tensile strain, we analyze the band widths of the relevant d-orbitals and the spin-split of the valence band around the K point, and the results are listed in Table S1. As can be seen, the spin-orbit coupling (SOC) induced spin-split is almost unchanged with an increase in strains, implying that the SOC effect may not be obviously influenced by the strains. This is unlike the case in the bent WSe2 nanoribbons, where the non-uniform bending strain induced flexoelectric effect can dramatically increase the SOC effect Tang et al. (2023).
The band widths of the d-orbital derived valence bands just below the Fermi level show a roughly unchanged trend with strains, and those of conduction bands just above the Fermi level show a slightly increasing trend with strains, as seen in Table S1. The relevant d-orbital band broadening is believed to be the main reason for the gap’s decreasing trend from 2H-MoS2, to 2H-MoSe2, and to 2H-MoTe2, with increasing atomic number of the chalcogen Wilson and Yoffe (1969). However, the total band width broadening of bands around the Fermi level cannot completely explain the band gap decrease of monolayer MoS2 with in-plane strains here. Comparing the case of strain $\epsilon_{x}=4\%$ with the non-strained one, G0W0 gives a gap decrease of 190 meV, see Table 1 in the main text, while the change of the total band broadening is only about 98 meV, see Table S1. So, the band gap decrease of monolayer MoS2 with strains should also involve the relevant band-center shifts.

### v3

@@ V Acknowledgements
The work of RKS was supported by the Department of Energy, Office of Science, Basic Energy Sciences, Materials Chemistry Materials Science and Engineering, under grant DE-DC0023356. The work of RKS and HT was partially supported by the Temple Center for Research on Energy Materials through the Catalytic Collaborative Funding Initiative sponsored by the Office of the Vice President for Research at Temple. The work of JPP was supported by DOE grant DE-SC0023356, and by the National Science Foundation, Division of Materials Research, Condensed Matter and Materials Theory, under grant DMR-2344734. We thank NERSC for computational resources. We thank Professor Maria Iavarone for helpful discussions and suggestions.
@@
In the 1H MoS2 monolayer, one Mo atom and its six adjacent S atoms form a trigonal prism (see Figure S1), and in the ligand field theory, the d-orbitals of the transition metal atoms split into three groups of non-bonding levels [5] The lower split group is mainly of $d_{z^{2}}$ character and is filled, and the fundamental gap is formed by this filled group and the unoccupied middle group.
To understand the trend in the change of gap with applied tensile strain, we analyze the band widths of the relevant d-orbitals and the spin-split of the valence band around the K point, and the results are listed in Table S1. As can be seen, the spin-orbit coupling (SOC) induced spin-split is almost unchanged with an increase in strains, implying that the SOC effect may not be obviously influenced by the strains. This is unlike the case in the bent WSe2 nanoribbons, where the non-uniform bending strain induced flexoelectric effect can dramatically increase the SOC effect [30]
The band widths of the d-orbital derived valence bands just below the Fermi level show a roughly unchanged trend with strains, and those of conduction bands just above the Fermi level show a slightly increasing trend with strains, as seen in Table S1. The relevant d-orbital band broadening is believed to be the main reason for the gap’s decreasing trend from 2H-MoS2, to 2H-MoSe2, and to 2H-MoTe2, with increasing atomic number of the chalcogen [31] However, the total band width broadening of bands around the Fermi level cannot completely explain the band gap decrease of monolayer MoS2 with in-plane strains here. Comparing the case of strain $\epsilon_{x}=4\%$ with the non-strained one, G0W0 gives a gap decrease of 190 meV, see Table 1 in the main text, while the change of the total band broadening is only about 98 meV, see Table S1. So, the band gap decrease of monolayer MoS2 with strains should also involve the relevant band-center shifts.
@@
GitHub repository for Input Files
You can download the input files from this link: https://github.com/Rajkishor46/MoS2Inputs.